OpenStudy (anonymous):

Find the solutions of g(x). Show each step. g(x) =2x^2 + 2x - 3

OpenStudy (anonymous):

I know there's two irrational solutions, that's about it

OpenStudy (anonymous):

g(x)=x^2+x-3/2 g(x)=x^2+x+1/4-1/4-3/2 g(x)=(x+1/2)^2-7/4

OpenStudy (anonymous):

It has three?

OpenStudy (anonymous):

Can you show me how you found those answers?

OpenStudy (anonymous):

i think we should solve g(x)=0 then we have (x+1/2)^2-7/4=0 (x+1/2)^2=7/4 x+1/2=7/4 --> x=5/4 x+1/2=-7/4 -->x=-9/4

OpenStudy (anonymous):

Thanks, and sorry I was eating lunch